COMMISSION 6 : ASTRONOMICAL TELEGRAMS (TELEGRAMMES ASTRONOMIQUES)

Report of Meeting, 23 August 1994

President: J. E. Grindlay Secretary: B. G. Marsden

President J. E. Grindlay, whose delayed overnight flight from Boston caused him to reach the assigned room in the Congresgebouw about five minutes after the meeting was due to start, opened the meeting by nominating R. M. West as incoming President for 1994–1997 and B. G. Marsden as incoming Vice-President. In accepting the latter nomination, Marsden remarked that, since Commission 6 provides oversight for the Central Telegram Bureau, it would be improper for him to accede to the position of President if he were at that time still Director of the Central Telegram Bureau. Members present then elected these officers and approved Grindlay and former President E. Roemer as members of the Organizing Committee. S. Isobe resigned from the Organizing Committee and was replaced by S. Nakano. Now that he has retired as Director of the National Astronomical Observatory of Japan, Y. Kozai felt that he should retire from the Commission, and he suggested that T. Nakamura should take his place. Grindlay suggested M. M. Phillips as a new member of the Commission. It was also suggested that there should be a representative from the Institute for Theoretical Astronomy in St. Petersburg, and G. R. Kastel’ was therefore proposed as a Consultant. D. W. E. Green, Associate Director of the Central Telegram Bureau, was reconfirmed as a Consultant.

In reporting on activities at the Central Telegram Bureau since the compilation of the report in IAU Trans. XXIIA Marsden noted that the sudden decision by the Smithsonian Institution during the second week of September 1993 no longer to absorb postage costs effective the beginning of October had made it necessary for the Bureau to take quick steps to avoid financial calamity. Those steps included an increase in the line charges (to $17.50 per line and $45.00 per item) and in the subscription rates for the printed Circulars (to $12.50 per month for regular accounts and $7.50 per month for special accounts—i.e., the same amounts as for the electronic version), the cessation of the requirement that subscribers to the electronic version of the Circulars also subscribe to the printed version, and—in collaboration with the Minor Planet Center—the transfer of urgent material on unusual minor planets to a new series of Minor Planet Electronic Circulars, which would not be made available in printed form and which initially used the same subscription list as the electronic version of the IAU Circulars. These changes, as well as the fact that the Bureau’s activity during the remainder of 1993 happened to be at a rather low level, staved off immediate disaster. Between August 1993 and August 1994 the number of subscribers to the printed Circulars decreased from 722 to 560, while the number of electronic subscribers increased from 249 to 342. Negotiations for alternative support from the Smithsonian—perhaps a waiving of the overhead charges levied on the Bureau—were continuing, but progress was very slow.

There followed some discussion, generally favorable, of the commission-restructuring plan proposed by the Executive Committee. Since it was proposed that Commission 6 remain directly under the authority of the Executive Committee, the effect of the plan on the Commission is expected to be minor at the present time.

Referring to the resolution adopted by Commission 6 in 1991, Grindlay noted with satisfaction that the Executive Committee’s budget for 1995–1997, printed in IAU Inf. Bull. No. 73, contained the sum of 11 700 CHF to help support the Central Bureau.

Commission 6 supported resolutions by Commissions 27 and 28 to the effect that the positions of novae and supernovae should be specified in the IAU Circulars in the J2000.0 system beginning 1995 Jan. 1.

Commission 6 also endorsed the resolution by Commission 20 concerning the major change to be introduced in the designation system for comets at the beginning of 1995. While no real change was being made in the naming system for comets at this time, it was noted that comets would not in the future be named immediately, but upon consultation between the Central Bureau and the Small Bodies Names Committee of Commission 20. This change was felt to be desirable, and in most instances it was expected that the naming process would be complete by the time preliminary orbit determinations were available.

The meeting concluded with some remarks by West and a general discussion on how new technological developments could further alter the manner in which the IAU Circulars were disseminated in the future. The recent experience of the collision between comet 1993e and Jupiter had shown the value—and the spirit of excitement—of real-time, world-wide communication among observers and the immediate availability of CCD images via the internet. Nevertheless, the problem existed that, while many of the images do largely speak for themselves, there could be no verification of the reliability of much of the information relayed. With the proliferation of electronic bulletin-boards and e-mail “exploders” the verification problem is becoming quite severe, and astronomers are often led on time-wasting wild-goose chases. The Central Bureau is renowned for its generally effective verification and rapid refereeing process, and it rarely disseminates information that is totally invalid. It is very desirable that this vetting should continue, and the IAU Circulars are a valuable source of initial information on important astronomical discoveries and can be appropriately cited in subsequent papers. But it would also be desirable to have the Circulars freely available electronically on WWW, the “world-wide web”. Since the IAU is itself in the process of establishing its own WWW server, it would seem very appropriate that it should contain the IAU Circulars as a matter of course. The sticking point is the need to support personnel to carry out the Bureau’s work—at all hours of the day and often at weekends and on holidays. Bureau Director Marsden is paid directly by the Smithsonian Institution, but Associate Director Green and other part-time support staff are not. Their salaries currently comes mainly from subscription charges. Ironically, if there were no subscription charges for the Circulars, the Smithsonian Institution would agree to pay the postal charges! Even without subscription charges, some administrative support would be necessary to maintain address lists for both the printed and electronic distribution of the Circulars. Despite the popularity of the WWW, many users prefer to receive the Circulars passively by e-mail, and some users still depend on seeing the printed Circulars, even though the latter are generally now delayed until five or even ten issues can be mailed at once. As new President, West resolved speedily to examine ways in which the Bureau’s staff could be assured financial support while the Circulars would be “freely” available.